# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

### v1

@@ 3.2.2 Predictive covariances using simulation
[…]
@@
7:$\hat{\Omega}_{p}\leftarrow Z_{pp}\Sigma_{p}Z_{pp}^{T}+\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(4)}\left(z_{i}^{(4)}\right)^{T}$
@@
[…]
@@
7:$\hat{\Omega}_{p}\leftarrow Z_{po}\Sigma Z_{po}^{T}+Z_{pp}\Sigma_{p}Z_{pp}^{T}-\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(4)}\left(z_{i}^{(4)}\right)^{T}$
Moreover, we can apply the Woodbury identity in (11) and obtain
@@
We have additionally considered an algorithm, where only the last term in (18) is approximated by simulation and all other terms in (11) are calculated deterministically. However, we observe in experiments that such an algorithm leads to less accurate estimates than when using Algorithm 2 and Algorithm 3 (results not tabulated).
@@ 3.2.3 Predictive variances using preconditioned Lanczos algorithms
Pleiss et al. (2018) use the Lanczos algorithm to approximate predictive covariance matrices for Gaussian process regression. We adapt and extend this approach for GMMs. Specifically, we approximate $\text{diag}(Z_{po}\Sigma Z^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T})$ in (18) by running the Lanczos algorithm with the matrix $\Sigma^{-1}+Z^{T}WZ$ using the normalized average of the column vectors of $Z^{T}WZ\Sigma Z_{po}^{T}$ as initial value and approximate $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T}\approx\tilde{Q}\tilde{T}^{-1}\tilde{Q}^{T}Z^{T}WZ\Sigma Z_{po}^{T}$, where $\tilde{Q}\tilde{T}\tilde{Q}^{T}\approx\Sigma^{-1}+Z^{T}WZ$ is obtained after $k$ steps of the Lanczos algorithm, $\tilde{Q}\in\mathbb{R}^{n\times k}$ has orthonormal columns, and $\tilde{T}\in\mathbb{R}^{k\times k}$ is tridiagonal. $\text{diag}(Z_{po}\Sigma Z^{T}WZ\Sigma Z_{po}^{T})$ and the other terms in (11) are deterministically computed, and predictive variances are thus approximated by
@@
The computation of a rank $k$ Lanczos approximation requires one matrix-vector multiplication with $\Sigma^{-1}+Z^{T}WZ$ in each iteration. Since the Lanczos algorithm suffers from numerical stability issues due to loss of orthogonality, we use a full reorthogonalization scheme (Wu and Simon, 2000). Computing the Lanczos approximation then has complexity $O(k^{2}m)$. Furthermore, we consider a similar way of preconditioning as in Kündig and Sigrist (2024) by applying the partial Lanczos algorithm to $P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}}\approx\tilde{Q}\tilde{T}\tilde{Q}$ and correcting for this by using $P^{-\frac{T}{2}}\tilde{Q}\tilde{T}^{-1}\tilde{Q}P^{-\frac{1}{2}}Z^{T}WZ\Sigma Z_{po}^{T}\approx(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T}$ as approximation. Note that the Lanczos algorithm requires a symmetric matrix and in contrast to the preconditioned CG method, we need to explicitly calculate a factor $P^{-\frac{1}{2}}$, where $P=P^{\frac{1}{2}}P^{\frac{T}{2}}$. This is possible for the diagonal, SSOR, and ZIC preconditioners. However, in our experiments, this form of preconditioning does not lead to more accurate predictive variance estimates (results not tabulated).
@@ 3.3 Software implementation
The methods presented in this article are implemented in the GPBoost library written in C++ with Python and R interfaces, see https://github.com/fabsig/GPBoost.***Krylov subspace methods can be enabled via the parameter matrix_inversion_method = "iterative" and the preconditioner is chosen via the parameter cg_preconditioner_type. For linear algebra calculations, we use the Eigen library version 3.4.99 and its sparse matrix algebra operations whenever possible. Multi-processor parallelization is done using OpenMP.
@@ 4 Convergence analysis
In the following, we theoretically analyze the properties of the Krylov subspace methods presented in this article. We denote by $\lambda_{\min}=\lambda_{m}\leq\dots\leq\lambda_{1}=\lambda_{\max}$ the eigenvalues of a symmetric matrix $A\in\mathbb{R}^{m\times m}$. The eigenvalues of the preconditioned matrix $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ determine the convergence speed and accuracy of the CG and SLQ methods. However, there are important difference among the two methods. It is known that the CG method converges fast if the majority of the eigenvalues are clustered together and an effective condition number $\kappa_{m-l,k}=\frac{\lambda_{k}}{\lambda_{m-l}}$ (Van der Sluis and van der Vorst, 1986) is small for small $k,l\geq 1$, despite the presence of a few outlying small and large eigenvalues; see, e.g., Van der Vorst (2003) and Nishimura and Suchard (2022). In particular, a large condition number $\kappa=\frac{\lambda_{\max}}{\lambda_{\min}}$ is not necessarily a problem for the convergence speed of the CG method. Pandolfi et al. (2024) have recently shown that, for the diagonal preconditioner and a Gaussian likelihood, the majority of the eigenvalues of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ are clustered together except for a few small eigenvalues. This likely explains the relatively popular application of the diagonal preconditioner for the CG method in the prior literature (Tsuruta et al., 2001; Taskinen et al., 2017; Garrick et al., 2019; Pandolfi et al., 2024). In line with this, our experiments in Section 5 show that the CG method, whether used with the diagonal, SSOR, or ZIC preconditioner, or even without preconditioning, yields similar runtimes.
@@
The situation is different for the SLQ method. Here, the smallest and largest eigenvalues, $\lambda_{\min}$ and $\lambda_{\max}$, are important for its convergence properties and accuracy as outlined in the following. In brief, the Lanczos algorithm factorizes a symmetric matrix as $A=QTQ^{T}$, where $Q\in\mathbb{R}^{n\times n}$ is orthonormal, and $T\in\mathbb{R}^{n\times n}$ is a tridiagonal matrix. This decomposition is computed iteratively, and $k$ iterations result in an approximation $A\approx\tilde{Q}\tilde{T}\tilde{Q}^{T}$, where $\tilde{Q}\in\mathbb{R}^{n\times k}$ contains the first $k$ columns of $Q$, and $\tilde{T}\in\mathbb{R}^{k\times k}$ the corresponding coefficients of $T$. Using $\log\det(A)=\operatorname*{tr}\log(A)$ and $\log(A)\approx\tilde{Q}\log(\tilde{T})\tilde{Q}^{T}$, SLQ then applies stochastic trace estimation to approximate $\log\det(A)$; see (15), Ubaru et al. (2017), and Kündig and Sigrist (2024, Appendix A.2) for more details. The logarithm function used in $\log(A)\approx\tilde{Q}\log(\tilde{T})\tilde{Q}^{T}$ is mostly determined by the extremal eigenvalues of $A=P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$. Furthermore, it is known that the Lanczos algorithm approximates the smallest eigenvalue $\lambda_{m}$ faster when the difference $\lambda_{1}-\lambda_{m}$ is smaller and when $\frac{\lambda_{m-1}-\lambda_{m}}{\lambda_{1}-\lambda_{m-1}}$ is larger (Golub and Van Loan, 2013). Similarly, the largest eigenvalue $\lambda_{1}$ is approximated the faster, the smaller $\lambda_{1}-\lambda_{m}$ and the larger $\frac{\lambda_{1}-\lambda_{2}}{\lambda_{2}-\lambda_{m}}$. Similar results can also be obtained for other eigenvalues. This means that the difference between the largest and smallest eigenvalues matters much more for the SLQ method compared to the CG method. In line with this, we find in our experiments in Section 5 that preconditioners for which the difference between the largest and smallest eigenvalues is small, such as the SSOR preconditioner, result in much more accurate log-determinant approximations compared to the diagonal preconditioner for which this difference is larger.
In addition to the above-mentioned intuitive and empirical arguments, there is convergence theory showing that the error of an SLQ-approximated log-determinant depends on the condition number $\kappa=\lambda_{1}/\lambda_{m}$ of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$. Below, we restate Theorem 3.2 of Kündig and Sigrist (2024) for our setting, which provides a stochastic error bound. We are not aware of any convergence results that relate an effective condition number $\kappa_{m-l,k}=\lambda_{k}/\lambda_{m-l}$ to the approximation error of an SLQ-approximated log-determinant.
@@
Let $\kappa=\frac{\lambda_{1}}{\lambda_{m}}$ denote the condition number of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$, $C_{mt}=\frac{Q_{\chi^{2}_{mt}}({1-\eta/2})}{mt}$, where $Q_{\chi^{2}_{mt}}(\cdot)$ is the quantile function of a $\chi^{2}-$distribution with $mt$ degrees of freedom, and $\widehat{\Gamma}=\frac{1}{t}\sum_{i=1}^{t}\|P^{-\frac{1}{2}}z_{i}\|_{2}^{2}e_{1}^{T}\log(\tilde{T}_{i})e_{1}$ an SLQ approximation. If the SLQ method is run with $l\geq\frac{\sqrt{3\kappa}}{4}\log\left(\frac{C_{mt}20\log(2(\kappa+1))\sqrt{2\kappa+1}}{\epsilon}\right)$ preconditioned CG steps and $t\geq\frac{32}{\epsilon^{2}}\log(\kappa+1)^{2}\log\left(\frac{4}{\eta}\right)$ number of random vectors, the following holds for any preconditioner $P$:
@@
In the following, we analyze the extremal eigenvalues of the preconditioned matrices $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for different preconditioners $P$. We denote by $\lambda_{\min}^{\text{P}}=\lambda_{m}^{\text{P}}\leq\dots\leq\lambda_{1}^{\text{P}}=\lambda_{\max}^{\text{P}}$ the eigenvalues of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for a preconditioner $P$. In addition, the eigenvalues of the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WZ$ are denoted by $\lambda_{\min}^{\text{none}}=\lambda_{m}^{\text{none}}\leq\dots\leq\lambda_{1}^{\text{none}}=\lambda_{\max}^{\text{none}}$
4.1 SSOR preconditioner
Theorem 4.2 and Corollary 4.1 derive the extremal eigenvalues, as well as lower and upper bounds, of the SSOR-preconditioned matrix $P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}$, for general non-Gaussian and Gaussian likelihoods, respectively. In particular, Theorem 4.2 and Corollary 4.1 imply that the largest and smallest eigenvalues, $\lambda_{\max}^{\text{SSOR}}$ and $\lambda_{\min}^{\text{SSOR}}$, of $P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}$ do not depend directly on neither the sample size $n$ nor the dimension of the random effects $m$ but rather only on the number of repeated observations per random effect, denoted by $d=\frac{n}{m_{1}}=\frac{n}{m_{2}}$ when having a balanced design. Furthermore, the results show that the ratio $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ and the difference between the largest and smallest eigenvalues of $P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}$ are increasing in the number of repeated observations per random effect; see, e.g., (26).
Theorem 4.2.
Denote $D_{1}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})\in\mathbb{R}^{m_{1}\times m_{1}}$ and $D_{2}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})\in\mathbb{R}^{(m-m_{1})\times(m-m_{1})}$. If $K=2$, the following hold true for the SSOR preconditioner defined in Section 3.1.1:
$P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}=I_{m}-\begin{pmatrix}0&0\\ 0&D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2}\end{pmatrix},$ (21)

### v2

@@ 3.2.2 Predictive covariances by simulating from multivariate Gaussian distributions
4:  Sample $z_{i}^{(1)}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,I_{m})$ and $z_{i}^{(2)}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,I_{n})$ and set $z_{i}^{(3)}\leftarrow\Sigma^{-\frac{1}{2}}z_{i}^{(1)}+Z^{T}W^{\frac{1}{2}}z_{i}^{(2)}$
@@
[…]
@@
4:  Sample $z_{i}^{(1)}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,I_{m})$ and $z_{i}^{(2)}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,I_{n})$ and set $z_{i}^{(3)}\leftarrow Z\Sigma^{\frac{1}{2}}z_{i}^{(1)}+W^{-\frac{1}{2}}z_{i}^{(2)}$
@@
[…]
We have also considered another version of such an algorithm. First, applying the Woodbury identity in (11), we obtain
@@
We can then approximate last term in (18) by simulation, and all other terms in (11) are calculated deterministically. However, we found that this algorithm leads to less accurate estimates than Algorithms 2 and 3 (results not shown).
@@ 3.2.3 Predictive variances using preconditioned Lanczos algorithms
pleiss2018constant use the Lanczos algorithm to approximate predictive covariance matrices for Gaussian process regression. We adapt and extend this approach for generalized mixed effects models. Specifically, we approximate $\text{diag}(Z_{po}\Sigma Z^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T})$ in (18) by running the Lanczos algorithm with the matrix $\Sigma^{-1}+Z^{T}WZ$ using the normalized average of the column vectors of $Z^{T}WZ\Sigma Z_{po}^{T}$ as initial value and approximate $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T}\approx\tilde{Q}\tilde{T}^{-1}\tilde{Q}^{T}Z^{T}WZ\Sigma Z_{po}^{T}$, where $\tilde{Q}\tilde{T}\tilde{Q}^{T}\approx\Sigma^{-1}+Z^{T}WZ$ is obtained after $k$ steps of the Lanczos algorithm, $\tilde{Q}\in\mathbb{R}^{n\times k}$ has orthonormal columns, and $\tilde{T}\in\mathbb{R}^{k\times k}$ is tridiagonal. $\text{diag}(Z_{po}\Sigma Z^{T}WZ\Sigma Z_{po}^{T})$ and the other terms in (11) are deterministically computed, and predictive variances are thus approximated as
@@
The computation of a rank $k$ Lanczos approximation requires one matrix-vector multiplication with $\Sigma^{-1}+Z^{T}WZ$ in each iteration. Since the Lanczos algorithm suffers from numerical stability issues due to loss of orthogonality, we use a full reorthogonalization scheme (wu2000thick). Computing the Lanczos approximation then has complexity $O(k^{2}m)$. Furthermore, we consider a similar way of preconditioning as in kundig2024iterative by applying the partial Lanczos algorithm to $P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}}\approx\tilde{Q}\tilde{T}\tilde{Q}$ and correcting for this by using $P^{-\frac{T}{2}}\tilde{Q}\tilde{T}^{-1}\tilde{Q}P^{-\frac{1}{2}}Z^{T}WZ\Sigma Z_{po}^{T}\approx(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T}$ as approximation. Note that the Lanczos algorithm requires a symmetric matrix and in contrast to the preconditioned CG method, we need to explicitly calculate a factor $P^{-\frac{1}{2}}$, where $P=P^{\frac{1}{2}}P^{\frac{T}{2}}$. This is possible for the diagonal, SSOR, and ZIC preconditioners. However, in our experiments, this form of preconditioning does not lead to more accurate predictive variance estimates (results not shown).
@@ 3.3 Software implementation
The methods presented in this article are implemented in the open-source C++ software library GPBoost, with accompanying high-level Python and R packages, see https://github.com/fabsig/GPBoost.***Krylov subspace methods can be enabled via the parameter matrix_inversion_method = "iterative" and the preconditioner is chosen via the parameter cg_preconditioner_type. For linear algebra calculations, we use the Eigen library version 3.4.99 and its sparse matrix algebra operations whenever possible. Multi-processor parallelization is done using OpenMP.
@@ 4 Convergence analysis
In the following, we theoretically analyze the properties of the Krylov subspace methods presented in this article. We denote by $\lambda_{\min}=\lambda_{m}\leq\dots\leq\lambda_{1}=\lambda_{\max}$ the eigenvalues of a symmetric matrix $A\in\mathbb{R}^{m\times m}$. The eigenvalues of the preconditioned matrix $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ determine the convergence speed and accuracy of the CG and SLQ methods. However, there are important differences among the two methods. It is known that the CG method converges fast if the majority of the eigenvalues are clustered together and an effective condition number $\kappa_{m-l,k}=\lambda_{k}/\lambda_{m-l}$ (van1986rate) is small for small $k,l\geq 1$, despite the presence of some outlying small and large eigenvalues; see, e.g., van2003iterative and nishimura2022prior. In particular, a large condition number $\kappa=\lambda_{1}/\lambda_{m}$ is not necessarily a problem for the convergence speed of the CG method. The situation is different for the SLQ method, for which the extremal eigenvalues are important for its convergence properties and accuracy, as described in the following.
@@
The Lanczos algorithm factorizes a symmetric matrix as $A=QTQ^{T}$, where $Q\in\mathbb{R}^{m\times m}$ is orthonormal, and $T\in\mathbb{R}^{m\times m}$ is a tridiagonal matrix. This decomposition is computed iteratively, and $k$ iterations result in an approximation $A\approx\tilde{Q}\tilde{T}\tilde{Q}^{T}$, where $\tilde{Q}\in\mathbb{R}^{m\times k}$ contains the first $k$ columns of $Q$, and $\tilde{T}\in\mathbb{R}^{k\times k}$ the corresponding coefficients of $T$. In brief, SLQ is based on $\log\det(A)=\operatorname*{tr}\log(A)$, $\log(A)\approx\tilde{Q}\log(\tilde{T})\tilde{Q}^{T}$, and stochastic trace estimation. The logarithm function used in $\log(A)\approx\tilde{Q}\log(\tilde{T})\tilde{Q}^{T}$ is mostly determined by the extremal eigenvalues of $A=P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$. It is known that the Lanczos algorithm approximates the smallest and largest eigenvalues $\lambda_{m}$ and $\lambda_{1}$ the faster, the smaller $\lambda_{1}-\lambda_{m}$ and the larger $(\lambda_{m-1}-\lambda_{m})/(\lambda_{1}-\lambda_{m-1})$ and $(\lambda_{1}-\lambda_{2})/(\lambda_{2}-\lambda_{m})$, respectively (golub2013matrix, Theorem 10.1.2, Corollary 10.1.3). Similar results can also be obtained for other eigenvalues. Moreover, even if the approximation due to the Lanczos algorithm was exact, stochastic trace estimators exhibit increased variance when the spectrum of the matrix $\log(A)$ is poorly conditioned. The extremal eigenvalues $\lambda_{m}$ and $\lambda_{1}$ are thus expected to be important for the accuracy of an SLQ-approximated log-determinant. Besides these heuristic arguments, there is convergence theory for SLQ-approximated log-determinants with stochastic error bounds that depend on the condition number $\kappa=\lambda_{1}/\lambda_{m}$ of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$. Below, we restate Theorem 3.2 of kundig2024iterative for mixed effects models.
@@
Let $\kappa=\lambda_{1}/\lambda_{m}$ denote the condition number of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$, $C_{mt}=\frac{1}{mt}Q_{\chi^{2}_{mt}}({1-\eta/2})$, where $Q_{\chi^{2}_{mt}}(\cdot)$ is the quantile function of a $\chi^{2}$-distribution with $mt$ degrees of freedom, and $\widehat{\Gamma}=\frac{1}{t}\sum_{i=1}^{t}\|P^{-\frac{1}{2}}z_{i}\|_{2}^{2}e_{1}^{T}\log(\tilde{T}_{i})e_{1}$ an SLQ approximation. If the SLQ method is run with $l\geq\frac{\sqrt{3\kappa}}{4}\log\left(\frac{C_{mt}20\log(2(\kappa+1))\sqrt{2\kappa+1}}{\epsilon}\right)$ preconditioned CG steps and $t\geq\frac{32}{\epsilon^{2}}\log(\kappa+1)^{2}\log\left(\frac{4}{\eta}\right)$ number of random vectors, the following holds for any preconditioner $P$:
@@
4.1 Extremal eigenvalues and convergence of the SLQ method
To understand the convergence properties of the SLQ method and to apply Theorem 4.1, we next analyze the extremal eigenvalues $\lambda_{\min}$ and $\lambda_{\max}$ and the condition numbers $\lambda_{\max}/\lambda_{\min}$ of the preconditioned matrices $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for the SSOR and diagonal preconditioners as well as the unpreconditioned matrix $\Sigma^{-1}+Z^{T}WZ$. We use the following notation.
Notation.
We denote by
$\lambda_{\min}^{\text{P}}=\lambda_{m}^{\text{P}}\leq\dots\leq\lambda_{1}^{\text{P}}=\lambda_{\max}^{\text{P}}$
the eigenvalues of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for a preconditioner $P$. The eigenvalues of the unpreconditioned matrix $\Sigma^{-1}+Z^{T}WZ$ are denoted by $\lambda_{\min}^{\text{none}}=\lambda_{m}^{\text{none}}\leq\dots\leq\lambda_{1}^{\text{none}}=\lambda_{\max}^{\text{none}}$. Furthermore, let
$d_{\max}=\max(Z^{T}Z)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ d_{\min}=\min(Z^{T}Z)$
denote the largest and smallest number of occurrences per random effect, and
$\sigma^{2}_{\min}=\min(\sigma^{2}_{1},\dots,\sigma^{2}_{K})\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sigma^{2}_{\max}=\max(\sigma^{2}_{1},\dots,\sigma^{2}_{K}).$
We call a random effects design “balanced” if every random effect $(b_{k})_{j}\in\mathbb{R}$, $j=1,\dots,m_{k}$, $k=1,\dots K$, occurs exactly $d$ times, i.e., $d=n/m_{k}=\sum_{i=1}^{n}(Z_{k})_{ij}$ for all $1\leq j\leq m_{k}$ and $k=1,\dots,K$.
In Lemmas 4.1 and 4.2, we first derive bounds for the extremal eigenvalues in the general case including Gaussian and non-Gaussian likelihoods. Proofs of Lemmas 4.1 and 4.2 can be found in Appendix A.1. Based on these results, we derive in Theorem 4.2 upper and lower bounds for the condition numbers if $K=2$ and the likelihood is Gaussian. If the random effects design is balanced, the condition numbers for the SSOR and diagonal preconditioners are known exactly. Theorem 4.2 also analyzes and compares theses condition numbers asymptotically for large $d_{\max}$ and $d$, i.e., when the matrix $\Sigma^{-1}+Z^{T}WZ$ is usually less sparse and preconditioners are thus even more important given that computations per CG iteration are more expensive. Furthermore, in Theorem 4.3, we derive upper bounds for the condition numbers for Bernoulli likelihoods.
Lemma 4.1.
If $K=2$, the following hold true for the SSOR preconditioner: